# Machine learning classification of black holes in the mass–spin diagram

### v1

@@ Abstract
We present the mass–spin diagram for classifying black holes and studying their formation pathways providing an analogue to the Hertzsprung–Russell diagram. This allows for black hole evolutionary tracks as a function of redshift, combining formation, accretion, and merger histories for the variety of black hole populations. A realistic black hole continuum constructed from initial mass and spin functions and approximate redshift evolution reveals possible black hole main sequences, such as sustained coherent accretion through cosmic time, i.e. Cosmic Accretion, or hierarchical merger trees. In the stellar-mass regime, we use a binary population synthesis software to compare three spin prescriptions for tidal evolution of Wolf-Rayet progenitors, showing how the mass–spin diagram exposes interesting modeling differences. We then classify black hole populations by applying supervised and unsupervised machine learning clustering methods to mass–spin datasets. While bare unsupervised clustering can nearly recover canonical population boundaries (stellar-mass, intermediate-mass, and supermassive), a more sophisticated approach utilizing deep learning via variational autoencoders for latent space representation learning aids in clustering of realistic datasets with subclasses that highly overlap in mass–spin space. We find that a supervised random forest can accurately recover the correct clusters from the learned latent space representation depending on the complexity of the underlying dataset, semi-supervised methods show potential for further development, and the performance of unsupervised classifiers is a great challenge. Our findings motivate future machine learning applications and demonstrate that the mass–spin diagram can be used to connect gravitational-wave and electromagnetic observations with theoretical models.
@@ I Introduction
Figure 1 displays an HR diagram from a synthetic stellar population generated with the COMPAS rapid isolated-binary evolution code with default parameter settings [6]. The stellar classifications, as the HR diagram reveals, are intimately tied to the phases of stellar evolution through the stars’ lifespans. We filter this prototypical theoretical population to obtain a small randomized subset whose evolutionary stages are shown in Fig. 1. In chronological order, these intermediate-to-high mass stars begin core hydrogen fusion, i.e. the main sequence (red dashed lines) and move through a very brief Hertzsprung gap characterized by shell helium burning, which is followed by drastic radial expansion as core helium fusion ignites and constitutes the dominant component of the stellar luminosity in the giant branch (dot-dash purple lines). Then these stars lose their envelopes, either through binary mass transfer processes or stellar wind mass loss, revealing the core-helium burning phase of stellar evolution (dotted green lines) which leads imminently to gravitational collapse, possibly associated with a supernova explosion and compact object formation. A few surviving white dwarfs (gray dots) and neutron stars (blue stars) occupy their respective lower luminosity regions with the neutron stars effective temperature being much higher due differences in their supernovae mechanisms. This highlights the utility of the HR diagram for organizing the variety of stellar types and diagnosing stellar evolution models. Extensions of the HR diagram are still being found for varied purposes [7, 8, 9, 3, 10], underscoring its impact in modern astronomy. However, black holes could only be drawn on the HR diagram if embedded in an accretion disk or due to non-classical possibilities such as Hawking radiation.
@@
Motivated by this paradigm, we ask: can an analogous diagram be constructed for astrophysical black holes? We propose that, similar to the HR diagram, one can visually explore how black holes can be organized and classified. While black holes are historically challenging to observe, their known macroscopic properties of mass and spin from General Relativity offer natural axes for classification. With electromagnetic (EM) techniques such as X-ray spectroscopy and radio interferometry providing constraints on stellar-mass and supermassive black holes [11, 12, 13, 14] and with gravitational wave (GW) astronomy now delivering catalogs of stellar-mass black hole mergers [15, 16, 17], we can consider how patterns in black hole properties may emerge.
@@
Emergent gravitational-wave population studies suggest a mass-dependent spin distribution in merging binaries [24], and prior studies have explored black hole mass, spin, and redshift relationships in GW data [25, 26, 27]. An observed anti-correlation in the $\chi_{\rm eff}$$q$ plane [28, 29, 30], where $\chi_{\rm eff}$ is the aligned effective spin parameter [31, 32], was subsequently confirmed in more recent analyses of gravitational wave catalogs [33, 34]. Structures can be inferred from the spin distributions using data-driven models that capture unanticipated features without inserting systematic bias [35]. Investigations into the influence of stellar rotation on black hole mass–spin outcomes has long been studied and is model dependent, see e.g. [36, 37], which often involves implicitly projecting onto the mass–spin plane. These works significantly advance understanding of stellar-mass black hole populations, and motivate development of a tool for interpreting black hole mass and spin measurements.
@@
Figure 2: A sketch of the mass–spin diagram for astrophysical black holes with dimensionless spin angular momentum (i.e. $\chi=c|\mathbf{S}|/Gm^{2}$) on the vertical axis. Regions are colored, as indicated in the legend, according to a selection of formation channels of black holes in single, binary, or higher systems across the mass spectrum: binary stars in isolation, dense stellar clusters, canonical population III stars, and hierarchical galaxy mergers. In reality, these formation channels will have overlap and their prevalence will depend on redshift. The two green vertical patches with red hatch symbolize the commonly used boundaries between the stellar-mass, intermediate-mass, and supermassive regimes. Typical astrophysical processes that drive the mass and spin growth across the diagram are written in each regime, with small arrows indicating uncertainty in the region borders for mass and spin predictions of formation channels. The solid blue line extending from the stellar mass regime to the upper right corner of the diagram represents the direction of the Cosmic Accretion main sequence for these black hole evolutionary tracks.
@@ II Black hole mass-spin diagrams
Observationally, mass is the first and easiest parameter to constrain. Observations of gravitational waves from coalescences of stellar-mass binaries with the LIGO/Virgo/KAGRA network (LVK) are most sensitive to a combination of the binary masses known as the chirp mass $\mathcal{M}$, and can offer weak constraints on spin magnitudes and directions typically expressed through the effective aligned-spin parameter $\chi_{\rm eff}$ [32, 31] and effective precession parameter $\chi_{\rm p}$ [49, 50], but the spin magnitudes can be better measured for certain systems [51, 52, 53]. The binary black hole spin-orbit misalignments, also known as spin tilts or orientations, are expected to carry imprints of their progenitor evolution and hold great implications for general relativistic spin precession, but they are not yet well constrained from GW data [54, 55, 56].
@@
Black holes are observed confidently in two main mass regimes: stellar-mass black holes through electromagnetic methods, such as accreting binaries [12, 59] or astrometry, and through the terrestrial gravitational wave detector network LVK [34]; and supermassive black holes in galactic nuclei and quasars, observed via multi-wavelength spectroscopy, gas and stellar dynamics, and reverberation mapping [60, 61, 13, 62, 63, 59], and hopefully via gravitational waves with a future space-based detector such as LISA via binary or minor mergers [64]. The intermediate-mass regime is populated with many strong candidates [20, 21] but has not yet been confidently confirmed beyond the few hundred solar mass extraordinary events from LVK. While these LVK events confirm the existence of black holes with masses greater than 100 $\rm\,M_{\odot}$ [65], it remains to be seen how the rest of the gap, ie from $\sim$1000 to $\sim 10^{5}$, is filled.
@@
In Figure 2 we present a schematic mass–spin diagram summarizing some main theoretical possibilities and expected trajectories. Each shaded region corresponds to a specific black hole formation channel, such as isolated stellar binaries, stellar clusters, hierarchical galaxy mergers, or Population III stars, and spans a range of spin magnitudes and black hole masses predicted by each scenario. Horizontal and vertical transitions are overlaid to indicate astrophysical processes relevant for growing the black hole masses and/or spins. The processes for intermediate-mass black holes, which may be populated by multiple overlapping formation pathways, are expected to be a mixture of processes identified from the stellar (i.e. stellar collapse, spin inheritance, tidal spin-up of the progenitor, and accretion) and super mass (dynamical friction, binary loss-cone stellar scattering, disk-driven migration and accretion) regimes. The commonly used boundaries between the three mass regimes are shown for reference in green regions with red hatching. The small arrows colored by formation channel indicate uncertainties on the theoretical boundaries of the formation channel. Therefore, this diagram serves as a visual analog to the HR diagram for stars: a two-dimensional landscape in which different populations occupy distinct, physically motivated regions, and evolutionary tracks and snapshots can be mapped and classified.
@@
We quantitatively illustrate this in Figure 3 which shows the redshift-dependent trajectories of black holes across the mass–spin plane under different physical growth scenarios, representing stellar-mass black holes (blue circles), intermediate-mass black holes (green squares), and supermassive black holes (red triangles). Each line, i.e. trajectory, traces the evolution of an individual black hole from redshift $z=10$ to $z=0$, with markers indicating the initial position in mass–spin space. The stellar-mass black holes are born with a power-law mass distribution ($m\sim m^{-2.3}$ [77]) and uniform spin magnitudes [78, 79], and evolve under an approximation to the classic model of [80] where we assume $z_{i}\rightarrow z_{j}$ with $i<j$ defines a redshift timestep. In this approximate model, a black hole of mass $m$ accretes an amount of mass $\Delta m(z)=m(z_{j})-m(z_{i})=m_{\rm n}(1+1.5(1-\bar{z}))$ where $\bar{z}=z/10$ and $m_{\rm n}$ is the natal mass, and undergoes corresponding spin increase via $\Delta\chi(z)=\chi_{i}+(\chi_{\rm eq}-\chi_{i})\eta\Delta m(z)/m$ where $\chi_{\rm eq}=0.998$ and $\eta=0.3$ is an ad-hoc parameter that controls the efficiency of spin-up. Intermediate-mass black holes are drawn from a log-flat mass distribution and bimodal initial spin distribution [81], and undergo modest stochastic mass growth $\Delta m(z)=m_{\rm n}(1+0.4(1-\bar{z}))(1+\mathcal{N}(0,0.003))$ accompanied with conditional spin drift (i.e., aggressive upward spin drift if $\chi<0.6$ at a given $z$), loosely modeling the combined effect of hierarchical mergers and chaotic accretion [82, 20]. We assume supermassive black holes originate from a 60%/40% mixture of light and heavy seed populations (light seeds from a Schechter-like distribution with natal spins $\sim\mathcal{N}(0.5,0.2)$ and heavy seeds from a log-uniform distribution with natal spins from $\sim\mathcal{N}(0.8,0.1)$) [83, 84, 73, 85, 86, 87, 88], and grow through coherent thin-disk accretion via $\Delta m(z)=m_{\rm n}(1+3.0e^{-\bar{z}/1.5})$ and with efficient spin-up modeled similar to the stellar-mass case. The resulting diagram shows the imprints of these evolutionary pathways in the mass–spin space, which cluster in regions over time and reveal mass–spin gaps (e.g.’s in the top-left, middle, and bottom right portions of the diagram) analogous to current mass gaps such as the pair instability and neutron star-black hole gaps.
@@
Next, we examine a COMPAS stellar-mass black hole population, acquired from a public repository [95] [^2] COMPAS data on Zenodo which assumed all stars have the same metallicity $Z=0.001$. Figure 4 shows the resultant black hole masses and spins where the Wolf–Rayet progenitor spins are evolved with tidal synchronization models. These models all use the same underlying binary tidal evolution framework of [96] for Wolf–Rayet star tidal evolution in the case where the binary is composed of a black hole enforcing tides on its Wolf-Rayet companion in the isolated binary formation channel. But there are major differences in their implementation of that framework. Fig. 4 shows the resulting dimensionless spin magnitudes $\chi_{2}$ of the black holes that formed from the initially less massive star as a function of their masses $m_{2}$. The points are color-coded by the orbital separation $a_{\rm preSN2}$ prior to the supernova explosion of the secondary which determines whether the spin is tidally synchronized. We compare three models [^3] Two such models were already implemented into a COMPAS post-processing script, which we have extended for obtaining a tidally synchronized Wolf-Rayet star’s spin: Qin et al. (2018) [97] in triangles, Bavera et al. (2020) [98] in dots, and Steinle & Kesden (2021) [99] in x’s. The spin prescriptions are applied to the same underlying COMPAS binary population, i.e., the same masses and separations are used for each spin model, ensuring consistent initial conditions across models. All shown binaries are selected to be merging within a Hubble time, and have pre-SN2 separations less than $100\,R_{\odot}$ as the tidal synchronization timescale is longer for larger $a_{\rm preSN2}$. This highlights how differing assumptions about the final evolutionary phase of the progenitor binary are seen in the mass–spin diagram. The different synchronization condition results in different numbers of tidally synchronized Wolf-Rayet stars, and hence different numbers of points in the diagram, relative to the underlying total population: $\approx 45\%$, $\approx 35\%$, and $\approx 80\%$ for Qin et al. (2018) in triangles, Bavera et al. (2020) in dots, and Steinle & Kesden (2021), respectively. In Qin et al. (2018) and Bavera et al. (2020), WR star spins are assumed to be tidally synchronized with the orbital motion when the orbital period is less than 1 day, and in Steinle & Kesden (2021) when the orbital semi-major axis is less than or equal to the analytic expression derived in Steinle & Kesden (2021) for the semi-major axis at synchronization (Eq. (13) of [99]).
@@
Figure 4: The predictions of three models for the spin magnitude of the Wolf-Rayet star tidally synchronized with a black hole companion in the isolated formation channel. The masses $m_{2}$ and spin magnitudes $\chi_{2}$ correspond to those of the secondary black holes (i.e., formed from the initially less massive star) in $\approx$ 2,500 binaries (from the COMPAS population [95] in Fig. 1) with semi-major axis $a_{\rm preSN2}<100\leavevmode\nobreak\ {\rm R}_{\odot}$ prior to its formation in a supernova explosion. The three spin models assume a different tidal synchronization condition, and are indicated by triangles (Qin et al. (2018)), circles (Bavera et al. (2020)) and x’s (Steinle & Kesden (2021)). The models of Qin et al. (2018) and Bavera et al. (2020) and of Steinle & Kesden (2021) are numerical and analytical extensions of the model of [96], respectively.
@@
This comparison reveals several key features of the mass–spin distribution expected from isolated binary evolution. Binaries with tight orbits ($a_{\rm preSN2}\lesssim 10\,R_{\odot}$) tend to produce high-spin black holes ($\chi_{2}\gtrsim 0.5$), while wider systems result in lower spins due to the assumption that inefficient angular momentum transfer during loss of the progenitor star’s envelope will produce a low natal Wolf-Rayet spin. The different spin models produce distinct trends in the mass–spin plane, indicating that spin magnitude is highly sensitive to the adopted physics of tidal evolution. For example, the Steinle & Kesden (2021) model has smooth dependence on the binary separation modulated by dependence on the mass ratio, which is distinct from the other two models whose points cloister nontrivially due to projection effects, with the assumptions of Qin et al. (2018) serving as a middle-ground between those of Bavera et al. (2020) and Steinle & Kesden (2021).
While a detailed comparison between these tidal spin-up models is beyond the scope of this work, we note that the key difference between the models, and hence between how well they fill the mass–spin space, is the use of analytic approximations in the model of Steinle & Kesden (2021), which self-consistently captures the tidal evolution dependence on the Wolf-Rayet mass, the binary mass ratio, and $a_{\rm preSN2}$, compared to the analytic fits in the models of Qin et al. (2018) and Bavera et al. (2020) which carry projection effects from the extra parameters. Thus, the predictions from Steinle & Kesden (2021) fill the space more physically (in mass and separation) compared to those from Qin et al. (2018) and Bavera et al. (2020). This implies that the model of Steinle & Kesden (2021) is ripe for implementation in population synthesis codes; further, it also provides self-consistent, non-linear coupled evolution of the Wolf-Rayet spin magnitude and spin direction under tidal synchronization and alignment, respectively, and accounts for stellar wind mass-loss, which are all necessary for realistic comparisons with LVK observations.
@@
Another important difference between these models is the use of the (second order) tidal coefficient $E_{2}$ that appears in the expression for the synchronization timescale of the equilibrium tide whose complicated dependence on stellar structure and composition is difficult to capture; the models of Qin et al. (2018) and Bavera et al. (2020) (both based on precomputed grids from the MESA stellar evolution code [100] assuming the tide only acts on the radiative envelope) use the same prescription for $E_{2}$ as given in Qin et al. (2018), while Steinle & Kesden (2021) use the classic prescription given by [101] which is implemented in rapid binary population synthesis codes such as StarTrack [102] and COMPAS [6]. Interestingly, these model predictions can share significant overlap, suggesting a testing ground for machine learning classification as we will see in Subsection III.2.

### v2

@@ Abstract
We present the mass–spin diagram for classifying black holes and studying their formation pathways providing an analogue to the Hertzsprung–Russell diagram. This allows for black hole evolutionary tracks as a function of redshift, combining formation, accretion, and merger histories for the variety of black hole populations. A realistic black hole continuum constructed from initial mass and spin functions and approximate redshift evolution reveals possible black hole main sequences, such as sustained coherent accretion through cosmic time or hierarchical merger trees. In the stellar-mass regime, we use a binary population synthesis software to compare three spin prescriptions for tidal evolution of Wolf-Rayet progenitors, showing how the mass–spin diagram exposes interesting modeling differences. We then classify black hole populations by applying supervised and unsupervised machine learning clustering methods to mass–spin datasets. While bare unsupervised clustering can nearly recover canonical population boundaries (stellar-mass, intermediate-mass, and supermassive), a more sophisticated approach utilizing deep learning via variational autoencoders for latent space representation learning aids in clustering of realistic datasets with subclasses that highly overlap in mass–spin space. We find that a supervised random forest can accurately recover the correct clusters from the learned latent space representation depending on the complexity of the underlying dataset, semi-supervised methods show potential for further development, and the performance of unsupervised classifiers is a great challenge. Our findings motivate future machine learning applications and demonstrate that the mass–spin diagram can be used to connect gravitational-wave and electromagnetic observations with theoretical models.
@@ I Introduction
Figure 1 displays an HR diagram from a synthetic stellar population generated with the COMPAS rapid isolated-binary evolution code with default parameter settings [6]. The stellar classifications, as the HR diagram reveals, are intimately tied to the phases of stellar evolution through the stars’ lifespans. We filter this prototypical theoretical population to obtain a small randomized subset whose evolutionary stages are shown in Fig. 1. In chronological order, these intermediate-to-high mass stars begin core hydrogen fusion, i.e. the main sequence (red dashed lines) and move through a very brief Hertzsprung gap characterized by shell helium burning, which is followed by drastic radial expansion as core helium fusion ignites and constitutes the dominant component of the stellar luminosity in the giant branch (dot-dash purple lines). Then these stars lose their envelopes, either through binary mass transfer processes or stellar wind mass loss, revealing the core-helium burning phase of stellar evolution (dotted green lines) which leads imminently to gravitational collapse, possibly associated with a supernova explosion and compact object formation. A few surviving white dwarfs (gray dots) and neutron stars (blue stars) occupy their respective lower luminosity regions with the neutron stars effective temperature being much higher due to differences in their supernovae mechanisms. This highlights the utility of the HR diagram for organizing the variety of stellar types and diagnosing stellar evolution models. Extensions of the HR diagram are still being found for varied purposes [7, 8, 9, 3, 10], underscoring its impact in modern astronomy. However, black holes could only be drawn on the HR diagram if they are embedded in an accretion disk or due to non-classical possibilities such as Hawking radiation.
@@
Motivated by this paradigm, we ask: can an analogous diagram be constructed for astrophysical black holes? We propose that, similar to the HR diagram, one can visually explore how black holes can be organized and classified. While black holes are historically challenging to observe, their known macroscopic properties of mass and spin from General Relativity offer natural axes for classification. With electromagnetic (EM) techniques such as X-ray spectroscopy and radio interferometry providing constraints on stellar-mass and supermassive black holes [11, 12, 13, 14] and with gravitational-wave (GW) detections now delivering catalogs of stellar-mass black hole mergers [15, 16, 17], we can consider how patterns in black hole properties may emerge.
@@
Emergent gravitational-wave population studies suggest a mass-dependent spin distribution in merging binaries [24], and prior studies have explored black hole mass, spin, and redshift relationships in GW data [25, 26, 27]. An observed anti-correlation in the $\chi_{\rm eff}$$q$ plane [28, 29, 30], where $\chi_{\rm eff}$ is the aligned effective spin parameter [31, 32] and $q$ the mass ratio, was subsequently confirmed in more recent analyses of gravitational-wave catalogs and helps reveal subpopulations in the observed dataset associated with different formation scenarios [33, 34, 35]. Structures can be inferred from the spin distributions using data-driven models that capture unanticipated features without inserting systematic bias [36]. Investigations into the influence of stellar rotation on black hole mass–spin outcomes has long been studied and is model dependent [37, 38], which often involves implicitly projecting onto the mass–spin plane. These works significantly advance understanding of stellar-mass black hole populations, and motivate development of a tool for interpreting black hole mass and spin measurements.
@@
Figure 2: A sketch of the mass–spin diagram for astrophysical black holes with dimensionless spin angular momentum (i.e. $\chi=c|\mathbf{S}|/Gm^{2}$) on the vertical axis. Regions are colored, as indicated in the legend, according to a selection of formation channels of black holes in single, binary, or higher systems across the mass spectrum: binary stars in isolation, dense stellar clusters, canonical population III stars, and hierarchical galaxy mergers. In reality, these formation channels will have overlap and their prevalence will depend on redshift. The two green vertical patches with red hatch symbolize the commonly used boundaries between the stellar-mass, intermediate-mass, and supermassive regimes. Typical astrophysical processes that drive the mass and spin growth across the diagram are written in each regime, with small arrows indicating uncertainty in the region borders for mass and spin predictions of formation channels. The solid blue line extending from the stellar-mass regime to the upper right corner of the diagram represents the direction of a Cosmic Accretion main sequence.
@@ II Black hole mass-spin diagrams
Observationally, mass is the first and easiest parameter to constrain. Observations of gravitational waves from coalescences of stellar-mass binaries with the LIGO/Virgo/KAGRA network (LVK) [15, 16, 17] are most sensitive to a combination of the binary masses known as the chirp mass, and can offer weak constraints on spin magnitudes and directions typically expressed through the effective aligned-spin parameter $\chi_{\rm eff}$ [32, 31] and effective precession parameter $\chi_{\rm p}$ [50, 51], but the spin magnitudes can be better measured for certain systems [52, 53, 54]. The binary black hole spin-orbit misalignments, also known as spin tilts or orientations, are expected to carry imprints of their progenitor evolution and hold great implications for general relativistic spin precession, but they are not yet well constrained from GW data [55, 56, 57].
@@
Black holes are observed confidently in two main mass regimes: stellar-mass black holes through electromagnetic methods, such as accreting binaries [12, 60] or astrometry, and through the terrestrial gravitational-wave detector network LVK [34]; and supermassive black holes in galactic nuclei and quasars, observed via multi-wavelength spectroscopy, gas and stellar dynamics, and reverberation mapping [61, 62, 13, 63, 64, 60], and hopefully via gravitational waves with a future space-based detector such as LISA via binary or minor mergers [65]. The intermediate-mass regime is populated with many strong candidates [20, 21] but has not yet been confidently confirmed beyond the few hundred solar mass extraordinary events from LVK. While these LVK events confirm the existence of black holes with masses greater than 100 $\rm\,M_{\odot}$ [66], it remains to be seen how the rest of the gap, ie from $\sim$1000 to $\sim 10^{5}$, is filled.
@@
In Figure 2 we present a schematic mass–spin diagram summarizing some main theoretical possibilities and expected trajectories. Each shaded region corresponds to a specific black hole formation channel, such as isolated stellar binaries, stellar clusters, hierarchical galaxy mergers, or Population III stars, and spans a range of spin magnitudes and black hole masses predicted by each scenario. Horizontal and vertical transitions are overlaid to indicate astrophysical processes relevant for growing the black hole masses and/or spins. The processes for intermediate-mass black holes, which may be populated by multiple overlapping formation pathways, are expected to be a mixture of processes identified from the stellar (i.e. stellar collapse, spin inheritance, tidal spin-up of the progenitor, and accretion) and supermassive (dynamical friction, binary loss-cone stellar scattering, disk-driven migration and accretion) regimes [20, 21]. The commonly used boundaries between the three mass regimes are shown for reference in green regions with red hatching. The small arrows colored by formation channel indicate uncertainties on the theoretical boundaries of the formation channel. Therefore, this diagram serves as a visual analog to the HR diagram for stars: a two-dimensional landscape in which different populations occupy distinct, physically motivated regions, and evolutionary tracks and snapshots can be mapped and classified.
@@
We quantitatively illustrate this in Figure 3 which shows the redshift-dependent trajectories of black holes across the mass–spin plane under different physical growth scenarios, representing stellar-mass black holes (blue circles), intermediate-mass black holes (green squares), and supermassive black holes (red triangles). Each line, i.e. trajectory, traces the evolution of an individual black hole from redshift $z=10$ to $z=0$, with markers indicating the initial position in mass–spin space. The stellar-mass black holes are born with a power-law mass distribution ($m\sim m^{-2.3}$ [78]) and uniform spin magnitudes [79, 80], and evolve under an approximation to the classic model of [81] where we assume $z_{i}\rightarrow z_{j}$ with $i<j$ defines a redshift timestep. In this approximate model, a black hole of mass $m$ accretes an amount of mass $\Delta m(z)=m(z_{j})-m(z_{i})=m_{\rm n}(1+1.5(1-\bar{z}))$ where $\bar{z}=z/10$ and $m_{\rm n}$ is the natal mass, and undergoes corresponding spin increase via $\Delta\chi(z)=\chi_{i}+(\chi_{\rm eq}-\chi_{i})\eta\Delta m(z)/m$ where $\chi_{\rm eq}=0.998$ and $\eta=0.3$ is an ad-hoc parameter that controls the efficiency of spin-up. Intermediate-mass black holes are drawn from a log-flat mass distribution and bimodal initial spin distribution [82], and undergo modest stochastic mass growth $\Delta m(z)=m_{\rm n}(1+1.5(1-\bar{z}))(1+\mathcal{N}(0,0.003))$ accompanied with conditional spin drift (i.e., aggressive upward spin drift if $\chi<0.6$ at a given $z$), loosely modeling the combined effect of hierarchical mergers and chaotic accretion [83, 20]. We assume supermassive black holes originate from a 60%/40% mixture of light and heavy seed populations (light seeds from a Schechter-like distribution with natal spins $\sim\mathcal{N}(0.5,0.2)$ and heavy seeds from a log-uniform distribution with natal spins from $\sim\mathcal{N}(0.8,0.1)$) [84, 85, 74, 86, 87, 88, 89], and grow through coherent thin-disk accretion via $\Delta m(z)=m_{\rm n}(1+3.0e^{-\bar{z}/1.5})$ and with efficient spin-up modeled similar to the stellar-mass case. The resulting diagram shows the imprints of these evolutionary pathways in the mass–spin space, which cluster in regions over time and reveal mass–spin gaps (e.g.’s in the top-left, middle, and bottom right portions of the diagram) analogous to current mass gaps such as the pair instability and neutron star-black hole gaps.
@@
Next, we examine a COMPAS stellar-mass black hole population, acquired from a public repository [96] [^2] COMPAS data on Zenodo which assumed all stars have the same metallicity $Z=0.001$. Figure 4 shows a subpopulation of black hole masses and spins where the Wolf–Rayet progenitor spins are evolved with three tidal synchronization models. These models all use the same underlying binary tidal evolution framework of [97] for Wolf–Rayet star tidal evolution in the case where the binary is composed of a black hole enforcing tides on its Wolf-Rayet companion in the isolated binary formation channel. But there are major differences in their implementations of that framework. Fig. 4 shows the resulting dimensionless spin magnitudes $\chi_{2}$ of the black holes that formed from the initially less massive star as a function of their masses $m_{2}$. The points are color-coded by the orbital separation $a_{\rm preSN2}$ prior to the supernova explosion of the secondary. We compare three models [^3] Two such models were already implemented into a COMPAS post-processing script, which we have extended. for obtaining the spin of a Wolf-Rayet star under tidal evolution: Qin et al. (2018) [98] in triangles, Bavera et al. (2020) [99] in dots, and Steinle & Kesden (2021) [100] in x’s. The spin prescriptions are applied to the same underlying COMPAS binary population, i.e., the same masses and separations are used for each spin model, ensuring consistent initial conditions across models. All shown binaries are selected to be merging within a Hubble time, and have pre-SN2 separations less than $100\,R_{\odot}$ as the tidal synchronization timescale is longer for larger $a_{\rm preSN2}$. This highlights how differing assumptions about the final evolutionary phase of the progenitor binary are seen in the mass–spin diagram. The different assumptions result in different numbers of tidally synchronized Wolf-Rayet stars, and hence different numbers of points in the diagram, relative to the underlying total population: $\approx 50\%$, $\approx 60\%$, and $\approx 20\%$ for Qin et al. (2018), Bavera et al. (2020), and Steinle & Kesden (2021), respectively. For Bavera et al. (2020) and Steinle & Kesden (2021), we assume WR star spins are tidally synchronized with the orbital motion when the orbital period is less than 1 day (and in the case of Steinle & Kesden (2021) we find that using Eq. (13) of [100] is a highly conservative limit for tidal synchronization), and for Qin et al. (2018) following their prescription.
@@
Figure 4: The predictions of three models for the spin magnitude of the Wolf-Rayet star tidally synchronized with a black hole companion in the isolated formation channel. The masses $m_{2}$ and spin magnitudes $\chi_{2}$ correspond to those of the secondary black holes (i.e., formed from the initially less massive star) in binaries from the COMPAS population [96] with color corresponding to the semi-major axis $a_{\rm preSN2}$ prior to its formation in a supernova explosion. The three spin models are indicated by triangles (Qin et al. (2018)), circles (Bavera et al. (2020)) and x’s (Steinle & Kesden (2021)). The models of Qin et al. (2018) and Bavera et al. (2020) and of Steinle & Kesden (2021) are numerical and analytical extensions of the model of [97], respectively.
@@
This comparison reveals interesting features of the mass–spin distribution expected from isolated binary evolution modeling. Binaries with tight orbits ($a_{\rm preSN2}\lesssim 10\,R_{\odot}$) tend to produce high-spin black holes ($\chi_{2}\gtrsim 0.5$), while wider systems result in lower spins due to the assumption that inefficient angular momentum transfer during loss of the progenitor star’s envelope will produce a low natal Wolf-Rayet spin. The different spin models produce different trends in the mass–spin plane, indicating that spin magnitude is sensitive to the adopted physics of tidal evolution. For example, the Steinle & Kesden (2021) model has smooth dependence on the binary separation modulated by dependence on the mass ratio, and produces more maximally spinning black holes whereas the other two models produce comparatively more slow rotators ($\chi<0.1$).
@@
While a detailed comparison between these tidal models is beyond the scope of this work, we note that a key difference between the models is the use of analytic approximations in the model of Steinle & Kesden (2021) compared to the analytic fits in the models of Qin et al. (2018) and Bavera et al. (2020). The predictions from Steinle & Kesden (2021) fill the space more uniformly but are less accurate relative to the details of stellar evolution. For example, these models differ in the use of the (second order) tidal coefficient $E_{2}$ that appears in the expression for the synchronization timescale of the equilibrium tide whose complicated dependence on stellar structure and composition is difficult to capture. The models of Qin et al. (2018) and Bavera et al. (2020) (both based on precomputed grids from the MESA stellar evolution code [101] assuming the tide only acts on the radiative envelope) use the same prescription for $E_{2}$ as given in Qin et al. (2018), while Steinle & Kesden (2021) use the prescription given by [102] which is often implemented in rapid binary population synthesis codes such as StarTrack [103] and COMPAS [6]. Interestingly, these model predictions can share significant overlap, suggesting a testing ground for machine learning classification as we will see in Subsection III.2.